### v1

# Exploring Missing Modality in Multimodal Egocentric Datasets

Abstract
Multimodal video understanding is crucial for analyzing egocentric videos, where integrating multiple sensory signals significantly enhances action recognition and moment localization. However, practical applications often grapple with incomplete modalities due to factors like privacy concerns, efficiency demands, or hardware malfunctions. Addressing this, our study delves into the impact of missing modalities on egocentric action recognition, particularly within transformer-based models. We introduce a novel concept—Missing Modality Token (MMT)—to maintain performance even when modalities are absent, a strategy that proves effective in the Ego4D, Epic-Kitchens, and Epic-Sounds datasets. Our method mitigates the performance loss, reducing it from its original $\sim 30\%$ drop to only $\sim 10\%$ when half of the test set is modal-incomplete. Through extensive experimentation, we demonstrate the adaptability of MMT to different training scenarios and its superiority in handling missing modalities compared to current methods. Our research contributes a comprehensive analysis and an innovative approach, opening avenues for more resilient multimodal systems in real-world settings.
1 Introduction
Multimodal video understanding has been the de facto approach for analyzing egocentric videos. Recent works have shown that the complimentary multisensory signals in egocentric videos are superior for understanding actions [32, 20, 21, 22, 28] and localizing moments [38, 2, 36]. However, multimodal systems need to be practical for real-world applications that could suffer from the incompleteness of modality inputs due to privacy, efficiency, or simply device failures [25]. For example, when predicting in real-time using a wearable device, parts of the recordings might be scrapped to preserve the privacy of the bystanders/camera wearer [14]. Furthermore, using all sensors could be expensive for a wearable device, opting for cheaper modalities such as audio or IMU. Thus, studying the impact of missing modalities is crucial for realistic performance expectations.
Figure 1: Most commonly, we train the multimodal models on modal-complete data. These models fail when testing data has inputs with missing modalities (orange plots). Our proposed adaptation to missing modality significantly improves the performance across datasets (green plots). When all test inputs are modal-incomplete ($r_{test}=100\%$), we surpass unimodal performance (blue plot) by 5 points in Epic-Kitchens, and we are on par with the unimodal accuracy in Epic-Sounds.
Still, the current effort to study the impact of missing modalities in egocentric datasets is limited. Most methods assume that all modality inputs are complete at training and inference time. Recent works have studied the effect of missing modalities for different tasks varying from recommendation systems to emotion recognition [33, 39, 31, 41, 24, 35]. Most works related to missing modalities have studied the problem at test time only [39, 35, 43, 41, 33, 3], while just a handful studied missing modalities in both training and testing time [24, 25, 30]. Similar to our setting, Lee et al. [25] propose a strategy to learn prompts for pre-trained backbones to deal with missing modalities. However, they analyze their method for image and text datasets only; we implement our version for action recognition and use it as a baseline in Section 4. More recently, Gong et al. [12] proposed a benchmark for multimodal generalization, focusing on few-shot learning recognition while considering missing modalities. Though the latter work proposes an interesting benchmark that includes a zero-shot and few-shot setup, no works have diagnosed how recent transformer-based approaches perform when modalities are missing for the action recognition setting.
In this work, we study the problem of missing modalities in egocentric action recognition. First, we investigate how current transformer-based models are affected by incomplete modalities at test time. In Figure 1, we observe how the current state-of-the-art audiovisual recognition model, Multimodal Bottleneck Transformer (MBT) [32], trained on modal-complete inputs, suffers from a critical degradation in performance when the missing modality rate increases. The advantage of the multimodal backbone (orange) is lost when the missing modality rate in the test set exceeds $\sim 30\%$ (Epic-Sounds) and $\sim 70\%$ (Epic-Kitchens), at which point the unimodal model (blue) becomes a better alternative. To address this problem, we propose learning the missing modality ”template” during training to replace missing modalities at test time. We call this template the Missing Modality Token (MMT) and explain how to learn it in Section 3.4. Figure 1 (Epic-Kitchens) also shows how our approach (green) dramatically improves the test accuracy and stays at least 5% points above the unimodal performance even when the test-set is fully modal-incomplete.
We verify the effectiveness of our method in 3 egocentric datasets, including Ego4D [14], which has a full coverage of RGB video but less than $70\%$ of the videos have audio. We extensively analyze our proposed training strategies, showing how to train with MMT under different missing modality scenarios. Our experiments show that our simple yet effective approach proposes a strong solution to this problem. Our contributions are threefold: (1) We present a thorough study of the missing modality problem in egocentric action recognition by considering effectiveness, efficiency, and the impact of the fusion strategy. (2) We propose Missing Modality Token (MMT) to deal with missing modalities at train and test time for egocentric video understanding. We propose two training strategies, namely fixed-replace and random-replace. (3) We extensively study our method and find that it significantly improves the presented baselines. Our work brings insights and paves the way to build multimodal backbones robust to missing modalities.
2 Related Work
Missing Modality Modeling. Addressing missing modalities presents a notable challenge, explored through various strategies by researchers from different areas. From medical applications [1] to sentiment analysis [3], missing modalities are a long-standing problem in multimodal understanding. Similarly, Ma et al. [30] and Colombo et al. [3] investigate missing modalities within a Bayesian Meta-learning framework. Meanwhile, Tsai et al. [39] and Zhao et al. [43] address unexpected missing or noisy modalities at test time using generative models to recreate missing inputs. Neverova et al. [33] focus on multimodal gesture recognition, employing depth, audio, and video streams, and updating network parameters based on different modality combinations. Our approach takes inspiration from [33] but uses a simpler learning approach that does not require different modality loss combinations. We learn a Missing Modality Token (MMT), which serves as a template for missing modalities during training without requiring complex pipelines using generative models.
Other studies utilizing transformers [40], such as the work of Parthasarathy et al. [35], explore missing modalities at test time and propose training-time augmentations. Radevski et al. [37] apply multimodal distillation to enhance training for models using only RGB at test time. Ma et al. [31] develop strategies for optimal fusion layers and class tokens in the context of missing modalities, focusing on image-text datasets. Our research differs by demonstrating effectiveness across various fusion layers (Section 4.5), indicating that optimization of this parameter is not crucial. Lee et al. [25] proposes to learn to prompt large multimodal backbones for image and text classification when modalities are missing at train and test time. We adapt their method to our setting and show that ours is more practical and effective for dealing with missing modalities in egocentric videos. Lastly, Gong et al. [12] introduce a benchmark for handling missing modalities within the Ego4D dataset, tailored for few-shot classification  [^1] Code and data are not available. Though the previous approach proposes an interesting benchmark it introduces an extra layer of complexity by including few-shot and zero-shot recognition into the missing modality problem. Our work proposes to diagnose and study the problem in a simpler setting to understand the effect of missing modalities in standard egocentric video understanding.
Multimodal Egocentric Video Understanding. Egocentric perception faces distinct challenges compared to traditional video understanding benchmarks such as ActivityNet [7] and Kinetics [19]. The nature of how egocentric datasets are captured means that they usually feature strongly aligned and synchronized audiovisual signals. Key benchmarks in this field, including Epic-Kitchens [4], Epic-Sounds [18], and the more recent and extensive Ego4D [14], have demonstrated the importance of audiovisual learning for understanding egocentric videos due to the complementary nature of the audio and visual modalities [38, 21, 20]. These datasets have facilitated the creation of several audiovisual backbones tailored for video understanding. Xiao et al. [42] introduced a CNN-based dual-stream architecture, utilizing SlowFast networks for the visual component [8] and a separate stream for audio [22]. With the advent and adaptability of transformer architectures, several studies have treated different modalities as input tokens for a multimodal transformer encoder [26, 23, 27, 10]. However, self-attention mechanisms can become prohibitively expensive as the number of tokens increases. To address this, Nagrani et al. [32] offered an efficient modality fusion technique that avoids costly self-attention in their Multimodal-Bottleneck Transformer (MBT). We build atop MBT and propose a method to make it robust for missing modalities at train and test time.
3 Dealing with Missing Modalities
This section details the aspects we consider while addressing the missing modality problem. Namely, the scenarios and evaluation (3.1), the multimodal design and fusion (3.2), the possible naive solutions to the problem (3.3), and our proposed method (3.4).
3.1 Problem Statement, Setup, and Evaluation
Given the training and testing multimodal data samples, let us denote the missing modality rates in each set with $r_{train}$ and $r_{test}$, respectively. These rates are computed by dividing the number of modal-incomplete samples by the total number of samples. Note that we only consider the setting of one modality being missing in the dataset. Our work discusses the strategies of training a multimodal model under two scenarios: training-with-missing, i.e. some samples in the training set are modal-incomplete ($r_{train}\neq 0\%$), or training-with-complete i.e. all samples in the training set all modal-complete ($r_{train}=0\%$).
To observe the trained model’s behavior under different missing modality severity levels, we create several variants of the test set by manually removing the modality information from the samples until $r_{test}=100\%$. For non-adapted models, we replace missing modality inputs at test time with dummy (empty) tensors.
Following previous works [31, 30] when experimenting with fully modal-complete datasets ($r_{train}=0\%$ and $r_{test}=0\%$), we assume the modality with the best unimodal performance (dominant) to be incomplete at test-time (e.g., audio for Epic-Sounds). Unlike previous works [30, 31, 25], we also validate our adaptation strategies on a dataset with naturally incomplete modalities in train and test splits. We use two modalities commonly available in the egocentric video datasets: visual (RGB frames) and audio. We evaluate classification accuracy on egocentric action recognition datasets.
3.2 Efficient and Effective Multimodal Fusion
We deal with missing modalities while considering the methods proven to be the most effective for multimodal fusion. Previous transformer-based methods addressing missing modalities looked mostly at basic methods, such as early or mid-fusion with cross-modal self-attention, where all tokens are concatenated at the fusion layer. This does not scale well in videos due to the attention mechanism’s quadratic complexity (to the input size) [32]. Instead, we use the current state-of-the-art audiovisual fusion, MBT [32], which proved to be more efficient and effective. The bottleneck transformer in the MBT design allows the model to distill and propagate the most essential information across modalities where each modality performs self-attention only with a small number of learnable ”bottleneck” tokens. Such design is especially useful for information-dense (redundant) modalities like video.
Architecture Input dummy tokens Audio $r_{test}$ Accuracy
Self-attention
0% 45.3%
50% 32.7%
75% 26.7%
100% 20.5%
Bottleneck
0% 55.5%
50% 34.3%
75% 23.7%
100% 13.6%
Bottleneck
0% 55.5%
50% 38.7%
75% 30.2%
100% 21.7%
Video-only
0% - 100% 41.4%
Table 1: Bottleneck fusion vs. vanilla self-attention with early fusion in Epic-Sounds dataset. Bottleneck fusion is shown with 2 ways of manipulating missing inputs at test time: passing dummy tokens or removing the tokens. All shown models fall far behind the unimodal (video) performance.
Figure 2: Learning with MMT explained. Left: Given modal-incomplete data, it is still unclear how to effectively train a multimodal model and then predict with it. Right: We demonstrate how we perform training and testing with modal-incomplete data by using Missing Modality Token (MMT). MMT learns the best representation of missing inputs. At training time, we use both modal-incomplete and modal-complete samples to train MMT. With the latter, we randomly decide for which samples to replace the signal tokens with MMT. This strategy allows us to effectively represent the missing modality tokens at test time.
Fusion Layer. A key aspect of a multimodal fusion approach is the design of the fusion layer ($L_{f}$). $L_{f}$ is the layer at which the cross-modal interactions happen. We observe the performance of the original bottleneck model with modal-complete audiovisual inputs and train the model with different fusion layers. We show the test accuracy for these models in Figure 6 (marked as Non-Adapted) for the Epic-Kitchens and Epic-Sounds datasets. We find that the performance does not change significantly with different fusion layers when $r_{test}=0$. However, the fusion layer does make a difference when inputs are incomplete (e.g., 35% test accuracy with $L_{f}=0$ vs. 42% with $L_{f}=11$ in Epic-Sounds at $r_{test}=50\%$), as shown in Section 4.5. Overall, fusing earlier is preferred in Epic-Kitchens, but fusing later gives better results in Epic-Sounds. This outcome is consistent with the observations from the previous work [31]: the best fusion strategy is dataset-specific. While this observation might be intuitive, it is impractical when the model is very sensitive to the fusion layer, as searching for the best layer might be computationally expensive. Thus, we analyze the effect of the fusion layer when modalities are missing and show the effect of our approach.
3.3 Model’s intrinsic capabilities with missing modalities
Before jumping to any training adaptation strategy, we first explore the capability of the models trained without any specific adaptation to the missing modality.
Bottlenecks Condensing the Information. As mentioned above, MBT uses a small set of learnable fusion tokens to exchange information between the modalities. One might wonder if the ability of the bottleneck to ”condense” the information already plays some role in dealing with missing inputs. To investigate this, Table 1 compares training a vanilla self-attention model with $L_{f}=0$ (row 1) with the training MBT with the same fusion layer (row 2). Table 1 shows that vanilla self-attention performs poorly in the multimodal setup with complete modalities, losing 10 points, compared to the bottleneck fusion. Surprisingly, when the modality inputs are incomplete, vanilla self-attention performs better (20% vs 13.6% with $r_{test}=100\%$). This shows that while the bottlenecks are effective for multimodal fusion, they are sensitive to missing modalities and they must be adapted to address incomplete modalities.
Only Passing Complete Inputs at Test-Time. As an alternative to passing an empty tensor when the modality is missing, one could suggest simply discarding the tokens from the missing signal and only feeding the transformer with the non-missing modality. The previous alternative is only possible when using transformers, as they can take a variable number of input tokens. Table 1 shows that this strategy (row 3) performs better than just passing empty tensors (row 2). Nevertheless, the model fails to reach video-only performance (row 4), even with a lower missing rate $r_{test}=50\%$, and only reaches 21.7% when $r_{test}=100\%$. Thus, this adaptation is still far from optimal, and we seek better solutions.
3.4 Our Approach to Dealing with Missing Modalities
We suggest a simple and generic way to deal with missing modalities. Instead of passing dummy inputs or discarding the tokens of the missing inputs at test time, we propose to learn a ”template” for the missing inputs. We introduce a learnable Missing Modality Token (MMT), which is trained to represent the tokens of the missing modality by learning from the tokens of the non-missing modality. As shown in Figure 2 (left) and explained in the previous sections, it is unclear how to deal with inputs with missing modalities. To this end, we introduce the Missing Modality Token (MMT). Figure 2 (right) illustrates how the MMT replaces the token of missing modalities. Our solution is parameter efficient since MMT is shared across samples. However, we still add positional embeddings to each token to preserve the location information of each missing patch. As shown in Figure 2 the MMT is learned at training time by alternating complete-input tokens with MMT. We propose two strategies to train with MMT: fixed-replace and random-replace and discuss how to apply them in training-with-complete and training-with-missing scenarios. At test time, we replace the missing modality inputs with our learned MMT.
Strategy fixed-replace. This strategy assumes that there is a fixed subset of the training instances with a missing modality. MMT tries to learn the best representation of the missing modality using this fixed subset. This strategy can be straightforwardly applied for training-with-missing scenario, as $r_{train}$ of the training samples are modal-incomplete. In training-with-complete, we manually create this fixed subset of modal-incomplete instances by scraping the modality data in some instances for the whole training process. Essentially, we convert this scenario to training-with-missing by making $r_{train}>0$. We input MMTs to the transformer to represent the tokens of the scrapped inputs. Note that when training-with-missing we can still control $r_{train}$ by removing the modality from more instances atop the ones that are naturally missing. In this strategy, we completely discard these samples’ modality information during the whole training.
Strategy random-replace. Instead of dropping the modality in a fixed subset of training samples, we remove the modality for the random samples (keeping $r_{train}$ unchanged). For each training sample, with probability $p$, the ”missing” modality tokens will be replaced with MMT. Our intuition is that by picking the samples randomly, rather than using a fixed pool, the model can ”see” the same samples in both missing and complete modality scenarios. By doing so, the model can learn a better representation of the missing inputs. In training-with-missing, there is still a fixed set of modal-incomplete training samples. We input MMTs for these samples, too. Basically, in this scenario, fixed-replace is given, and we randomly drop from the modal-complete subset.
Inference. Regardless of the strategy used, we replace the missing inputs tokens with the learned MMT at test time.
4 Experiments
We present a detailed analysis of our MMT under both training-with-complete and training-with-missing. For both scenarios, we explore both strategies fixed-replace and random-replace to train MMT (Sections 4.34.4). We report classification accuracy under several missing rates $r_{test}$ for three egocentric datasets. Additionally, we study the effect of fusion layers $L_{f}$ in Section 4.5. Then, we compare our method vs. a baseline proposed by [25] in Section 4.6.
4.1 Datasets
We use videos from Ego4D [14] for pre-training the MBT backbone. Due to privacy and regulations, only 2.5K of 3.7K video hours have original audio in this dataset. We trim 450K 10-second modal-complete audiovisual clips spanning all 2.5K modal-complete hours. For the downstream tasks, we use the following egocentric action recognition benchmarks:
Epic-Kitchens-100 [5] has 90K trimmed clips of variable length, spanning 100 video hours. Each clip is labeled with a noun + verb pair, which describes the camera-wearer action. In total, there are 300 noun and 97 verb classes in the dataset. We train the model with 2 heads to jointly predict verb and noun classes. All videos in the dataset have complete visual and audio streams ($r_{train}=r_{test}=0\%$).
Epic-Sounds [18] spans the same 100 video hours as Epic-Kitchens but is annotated with sound labels. This dataset does not follow the noun and verb annotations from Epic-Kitchens, instead, it has 44 unique class labels. The dataset is composed of 79K annotated clips.
Ego4D-AR: We use the annotated clips from the Short-Term Action Anticipation task in Ego4D benchmark [14] to create an action recognition dataset that we dub as Ego4D-AR [^2] Ego4D does not have an action recognition benchmark. Specifically, we use the provided time-to-contact timestamps to trim the clips and the anticipated actions as labels. Ego4D-AR contains 142K clips annotated as noun and verb pairs. Overall, there are 128 noun and 81 verb classes. We find that the verb classes are highly imbalanced in this dataset. Therefore, we balance the class weights in the cross-entropy loss during training. We provide more details on the dataset in Supplementary. Similarly to Epic-Kitchens, we use 2 heads to predict the nouns and verbs. As we didn’t filter the Ego4D videos for this dataset, it has naturally missing modality. Only $71\%$ of training clips and $73\%$ of test clips have audio ($r_{train}=29\%$, $r_{test}=27\%$).
For clarity and due to space constraints, we report the verb accuracy for Epic-Kitchens, the class accuracy for Epic-Sounds, and the noun accuracy for Ego4D-AR in this section. We report the rest of the metrics for Supplementary.
Dataset Audio Video Audiovisual
Epic-Kitchens 40.0% 63.2% 64.0%
Epic-Sounds 46.5% 41.4% 55.2%
Ego4D-AR 26.3% 34.6% 36.4%
Table 2: The performance of the audio, video, and bottleneck audiovisual models on each dataset. We train and evaluate all models with $r_{train}$ = 0. In all datasets, multimodal performance beats unimodal performance.
Figure 3: Missing modality rate in training data $r_{train}$ vs. accuracy when training with fixed-replace strategy. We show that our MMT is effective in learning with missing modalities under several $r_{train}$ regimes. For Epic-Sounds and Ego4d-AR the higher masking seems to be more beneficial to learn MMT. However, for Epic-Kitchens lower masking yields better accuracy overall.
In Table 2, we show the unimodal and multimodal performance for each downstream dataset. As expected from the annotation strategy, video is the dominant modality in Epic-Kitchens and Ego4d-AR, and audio is the one in Epic-Sounds. Therefore, as mentioned in Section 3.1, we train our models to be robust to missing video in Epic-Kitchens and Ego4D-AR and to missing audio in Epic-Sounds. Since Epic-Sounds has fewer training samples and has a more balanced unimodal performance of each modality, we use it more extensively in our analysis.
For training-with-complete, we use Epic-Sounds and Epic-Kitchens. For training-with-missing, we use Ego4D-AR and create a modal-incomplete version of Epic-Sounds by enforcing $r_{train}>0$.
4.2 Implementations Details
Pre-training. We use the audiovisual MAEs [15, 9, 11] protocols and train our own implementation of Audiovisual Bottleneck MAE. We use the trimmed Ego4D clips and train for 200 epochs. We mask 70% of audio and 90% of video tokens. We use the same pre-trained model for all experiments. More details of the decoder used in the pre-training can be found in the supplementary material.
Architecture. Following [17], we use ViT-Base [6] with 12 transformer layers, 12 attention heads, and embedding dimension 768 as the encoder for each modality. For the fusion design, we follow MBT [32] and fix the number of bottlenecks to $B=4$ and the fusion layer to $L_{f}=8$ (except for the fusion layer ablation).
Inputs. Following [13, 17], we convert an audio waveform of $t$ seconds to log Mel-filterbank with 128 Mel-frequency bins, with a Hanning window of 25ms, shifting every 10ms. The output is a spectrogram of $128\times 100t$. We use 8-second audio and the patch size of $16\times 16$, resulting in $(128\times 100\times 8)/256=400$ audio tokens. For video, we sample 16 RGB frames at $8$ fps of $224\times 224$. Similarly to [17], we tokenize the frames with 3D convolutions, using the spacetime patch size of $16\times 16\times 2$. Each video input produces (16×224×224)/(256×2)=156816$\times$ 224$\times$ 224)/(256$\times$ 2)=156816 × 224 × 224 ) / ( 256 × 2 ) = 1568 tokens.
Finetuning. We train for 50 epochs in Epic-Kitchens experiments, 20 in Epic-Sounds, and 15 in Ego4D-AR. We use SpecAugment [34] for audio augmentation and Augmix [16] for video augmentation. We use AdamW [29] optimizer with half-cycle cosine learning rate decay.
Making training sets with different fixed $r_{train}$. We begin by taking all training instances that are complete across all modalities and shuffle them randomly once. For the Ego4D-AR dataset, instances that are inherently missing are placed at the very start of this sequence. Following this, we establish our desired missing rate, represented by $r_{train}$ by sampling from the start of this list. This method ensures that any increase in $r_{train}$ builds upon the existing set of modal-incomplete instances, meaning that if we increase $r_{train}$, we add new modal-incomplete instances to those already included, thereby maintaining a cumulative effect.
4.3 The effect of the number of modal-incomplete samples $r_{train}$ in fixed-replace.
We analyze the effect of $r_{train}$, the proportion of the modal-incomplete inputs in the training data, under the fixed-replace strategy in each dataset. For Epic-Kitchens we use $r_{train}\in\{25\%,50\%,75\%\}$, and $r_{train}\in\{29\%,50\%,75\%\}$ for Ego4D-AR. Recall that Ego4D-AR has a natural $r_{train}=29\%$, and we can’t reduce this value. We observed that higher values of $r_{train}$ yield better performance in Epic-Sounds; hence, we use $r_{train}\in\{40\%,60\%,80\%\}$ in this dataset.
Figure 3 shows the results for each dataset. We can see how Epic-Sounds and Ego4D-AR follow a similar trend: with the lowest $r_{train}$ in each dataset, the model does not adapt well to the missing inputs. For example, at $r_{test}=100\%$ in Ego4D-AR, the model trained with $r_{train}=50\%$ is $8\%$ points more accurate than the model trained with $r_{train}=29\%$. Furthermore, training with higher $r_{train}=50\%$ and $r_{train}=75\%$ still yields reasonable accuracy at lower $r_{test}$. The opposite is true for Epic-Kitchens, where not only does the model trained with the lowest $r_{train}=25\%$ adapt well to the missing modality, but the accuracy drops significantly when we train with more modal-incomplete inputs. Overall, we highlight that: (1) for all datasets, higher $r_{train}$ results in higher accuracy on the fully modal-incomplete test set ($r_{test}=100\%$), emphasizing the importance of training MMT with more samples; (2) surprisingly, while higher $r_{train}$ causes more modality information loss, we can still train accurate models in Epic-Sounds and Ego4D-AR.
Figure 4: Modality drop probability $p$ vs. accuracy for modal-complete Epic-Sounds and Epic-Kitchens. In Epic-Sounds, as $p$ decreases to $40\%$ the model’s performance with missing audio goes down. While dropping with high $p=80\%$, the model still reaches decent performance at all $r_{test}$. In Epic-Kitchens, we observe very similar behavior as in fixed-replace.
4.4 The effect of the Modality Drop Probability $p$ in random-replace
We analyze the impact of $p$, the probability with which the tokens of modal-complete inputs are replaced with MMT. For easier comparison with the results in 4.3, we use $p\in\{25\%,50\%,75\%\}$ for Epic-Kitchens and Ego4D-AR, and $p\in\{40\%,60\%,80\%\}$ for Epic-Sounds.
Training-with-complete Figure 4 shows changes in performance with different $p$ in training-with-complete. We observe how $p$ exhibits a similar effect in random-replace as $r_{train}$ in fixed-replace. However, there are some noticeable differences. First, across all $p$ values, the models trained with random-replace reach the same or higher accuracy compared to the corresponding models trained with fixed-replace (Figure 3). This is especially evident in Epic-Sounds: by switching from fixed-replace to random-replace, we get an average improvement of $2.3\%$, $1.6\%$, and $1.4\%$ points for $p=r_{train}$ of $40\%$, $60\%$, and $80\%$, respectively. We attribute the greater improvement with the lower $p=40\%$ to the ability of the model to ”see” more of the same samples appear in both missing and complete modality scenarios, as we outline in 3.4. When $p$ is too high, fewer samples have a chance of being seen by the model as modal-complete multiple times during training. Furthermore, while the improvement is the lowest for $p=80\%$, this model still performs optimally at $r_{test}=0\%$, which is not the case for the model with $r_{train}=80\%$. Thus, we find random-replace more effective for training MMT.
Training-with-missing In Figure 5, we show the results of training with random-replace in training-with-missing scenario. Note how we are using both $r_{train}$ and $p$ here. Because of that, the numbers are hard to compare to other results directly, but we can still see similar patterns. Consistently with all previous results, the more samples we replace with MMT during training, the better the model adapts to severely missing modality at test time. Furthermore, dropping the modality of too many inputs (randomly or fixed) negatively affects the performance. For example, while the model trained with $p=75\%$ achieves higher accuracy at $r_{test}=100\%$ in Ego4D-AR, it seems that it learns to ignore the audio completely, as the performance at $r_{test}=0\%$ drops to the unimodal video performance in this dataset. Additionally, by distilling the results in Figures 3 and 5, we observe that using random-replace in Epic-Sounds with $r_{train}>0$ is more effective for training with missing modalities.
Figure 5: Results with the modal-incomplete training data. As Epic-Sounds does not naturally have missing modality in the training data, we manually remove the audio from $r_{train}=25\%$ = and $r_{train}=50\%$ of samples in the train set. In Ego4D-AR, there is a naturally missing audio in $r_{train}=29\%$ of the samples.
Figure 6: Fusion Layer $L_{f}$ vs. accuracy in the models trained with no adaptation strategy (Non-Adapted) and trained with MMT (Adapted). For training with MMT, we use random-drop strategy with $p=60\%$ for Epic-Sounds and $p=25\%$ for Epic-Kitchens. This strategy makes MBT more robust to missing modalities across all $L_{f}$ and significantly reduces the negative effect of missing modalities.
4.5 The effect of the fusion layer.
As we mentioned in Sec. 3.2, the fusion layer does affect the performance of the bottleneck model, especially when test inputs are modal-incomplete. In Fig. 6, we examine whether training with Modality Masking Training (MMT) enhances robustness to a missing modality at various fusion layers and whether models trained with MMT (adapted) show the same sensitivity to the fusion layer as those trained without it (non-adapted). As we observe, the introduction of MMTs makes the model more robust to the missing modality across all fusion layers in both Epic-Sounds and Epic-Kitchens. With extremely severe $r_{test}=100\%$, the adapted models perform with $\sim 45\%$ accuracy in Epic-Kitchens, while the unimodal audio model achieves $40\%$ in this dataset. Furthermore, these adapted models do not exhibit similar sensitivity to the fusion layer as non-adapted. For example, in Epic-Sounds, the non-adapted models trained with $L_{f}=11$ exhibit superior performance, which is not the case for the adapted models (at $r_{test}=100\%$, $37\%$ accuracy with $L_{f}=11$ but $\sim 40\%$ for other $L_{f}$). Overall, all adapted models perform consistently well, each providing decent performance in a modal-incomplete inference. Our approach effectively addresses a long-standing issue of selecting the appropriate fusion layer when faced with missing modalities [31].
4.6 MMT vs. other methods
In Tables 6(a)6(b) and 6(c), we report the test results using our fixed-replace and random-replace strategies and the baseline method for Epic-Sounds, Epic-Kitchens, and Ego4D-AR, respectively. We also report the accuracy of the unimodal model in each dataset. For the baseline, we implement the multimodal prompts method [25]. It was originally based on ViLT [23] for image-text classification, so we implemented our version for action recognition based on MBT. Note that this method relies on a strong pre-trained backbone to efficiently finetune it by optimizing a few network parameters. As we deal with audiovisual learning in egocentric videos, we rely on large-scale egocentric pre-training using MAEs. As seen in Tables 6(a)6(b)6(c), our strategy outperform the alternative proposed by [25].
$r_{test}$ 0% 25% 50% 75% 100%
unimodal 41.4 41.4 41.4 41.4 41.4
no adaptation 55.2 45.6 37.1 28.3 19.5
mm prompts [25] 36.7 33.3 30.0 26.1 22.6
fixed-replace(ours) 55.3 51.2 47.6 43.4 40.2
random-replace(ours) 56.3 52.3 48.6 44.6 40.7
$r_{test}$ 0% 25% 50% 75% 100%
unimodal 40.0 40.0 40.0 40.0 40.0
no adaptation 63.9 55.5 46.8 37.9 29.5
mm prompts [25] 32.5 31.1 30.2 29.2 28.2
fixed-replace (ours) 63.4 58.8 54.1 49.5 44.8
random-replace (ours) 63.4 59.0 54.6 50.0 45.3
$r_{test}$ 27% 50% 75% 100%
unimodal 34.6 34.6 34.6 34.6
no adaptation 30.5 25.9 21.1 16.0
mm prompts [25] 18.4 16.4 13.9 11.5
fixed-replace (ours) 36.2 34.9 33.6 32.4
random-replace (ours) 36.7 35.2 33.5 32.0
Table 3: Comparison of our method with baselines on various datasets. We demonstrate the accuracy of our method vs. the adaptation of the method proposed by [25] and unimodal performance across different missing modality ratios $r_{test}$. We show in bold the best result and underline the runner-up.
We find that across all datasets, our fixed-replace and random-replace strategies either reach (in Epic-Sounds and Ego4D-AR) or exceed (in Epic-Kitchens) the unimodal performance in extreme $r_{test}=100\%$. Interestingly, in Epic-Sounds, random-replace also regularizes the training and increases the $r_{test}=0\%$ performance (Table 6(a)) by 1.1%.
5 Conclusion
We explore the missing modality problem in multimodal egocentric datasets. We suggest a simple yet effective method by learning the optimal token representation of the missing modality (MMT). Placing learnable tokens to represent missing inputs provides an easy and intuitive way to train and test with modal-incomplete inputs. We propose 2 ways to learn MMT when training action recognition models and show how their performance brings us closer to robust and effective multimodal systems.